OpenStudy (flvskidd):

Photon336 At the city museum, child admission is $5.70 and adult admission is $9.70 . On Monday, 163 tickets were sold for a total sales of $1277.10 . How many adult tickets were sold that day?

OpenStudy (flvskidd):

Can you help me with one more?

OpenStudy (photon336):

A = number of adult tickets C = number of child tickets You need to make two equations for this again. \[5.70(C)+9.70(A) =1277.10\] \[A+C = 163 \] then solve these two by substitution

OpenStudy (flvskidd):

gotcha one sec

OpenStudy (flvskidd):

I got 82.9 but I know it's wrong

OpenStudy (photon336):

yeah that would be wrong because remember # of tickets has to be a whole number.

OpenStudy (photon336):

163-A = C Solve the second equation for A. 5.70(163-A)+9.70(A) = 1277.10

OpenStudy (flvskidd):

I got 87 and it was right. Thank you!

OpenStudy (photon336):

yes that's it do you know why this is the answer?

OpenStudy (flvskidd):

why?